OpenStudy (anonymous):

Profits for a large corporation was 4×10^5 dollars in July and 6×10^5 dollars in August. What was their combined profit in July and august ? PLZZ HELP !

OpenStudy (anonymous):

\[\large 4.0 \times 10^5+6.0 \times 10^5=10.0 \times 10^5\]It would be better though to go ahead and move the decimal, so it would be \(\large 1.0 \times 10^6\)

OpenStudy (anonymous):

Thank you so much !

OpenStudy (anonymous):

@Cshrix

OpenStudy (anonymous):

You're welcome! :)

OpenStudy (anonymous):

Can you help me with another problem ?